OpenStudy (anonymous):

Find the fourier transform of\[e^{-ax^2}\] \[F[f(\xi)]=\int\limits_{-\infty}^{\infty}f(x)e^{-i \xi x}dx\] You will need to know this fact:\[2 \int\limits_{0}^{\infty}e^{-ax^2}dx=\sqrt{\frac{\pi}{a}}\] I can get to \[2\int\limits_{0}^{\infty}e^{-ax^2}\cos(\xi x)dx\]

OpenStudy (anonymous):

u have to calculate last integral?

OpenStudy (anonymous):

Yup

OpenStudy (anonymous):

Ans is attached btw.

OpenStudy (anonymous):

The next step has to do with differentiating the inegral: \[L(\xi)=2\int\limits_{0}^{\infty}e^{-ax^2}\cos(\xi x)dx\]\[L'(\xi)=-2\int\limits_{0}^{\infty}xe^{-ax^2}\sin(\xi x)dx\] But not really sure how that got differentiated.

OpenStudy (anonymous):

\[F(a,b)=\int\limits_{0}^{\infty}e^{-ax^2}\cos(b x)dx\] \[\frac{\partial F}{\partial b}=-\int\limits_{0}^{\infty}xe^{-ax^2}\sin(b x)dx\] now use integration by parts to show that \[\frac{\partial F}{\partial b}=-\frac{b}{2a} F\] solve for \(F\)

OpenStudy (anonymous):

I knew that hehe, brain fart

OpenStudy (anonymous):

lol...:)

OpenStudy (anonymous):

Made it through... such a strange problem, turned in a diffeq... gonna be a fun test :)